# Patch-Level Tokenization with CNN Encoders and Attention for Improved Transformer Time-Series Forecasting

### v2

@@ I Introduction
Accurately forecasting multivariate time-series data remains a challenging problem due to the complex, noisy, and often non-stationary nature of many real-world temporal processes. Such data are influenced by a combination of short-term fluctuations, longer-term trends, and interactions among multiple variables, resulting in non-linearity, heterogeneous feature dependencies, and evolving temporal dynamics. These characteristics complicate the design of effective forecasting models across a wide range of application domains.
Classical approaches to time-series forecasting, including autoregressive and moving-average models [12, 3], rely on assumptions of linearity and stationarity that are frequently violated in practice. As a result, deep learning methods have been increasingly adopted for modelling complex temporal behaviour. Recurrent neural networks (RNNs) [2], particularly Long Short-Term Memory (LSTM) [7] and Gated Recurrent Unit (GRU) [13] architectures, are among the most widely used models due to their ability to process sequential data and capture temporal dependencies. However, recurrent models summarize historical information through compressed hidden states, which can limit their ability to retain long-range dependencies and may lead to unstable training behaviour when long input sequences are considered.
@@
The proposed method is evaluated on synthetic multivariate time-series datasets designed to reflect key structural properties commonly observed in real-world temporal data, including non-linear dynamics, interactions between static and dynamic factors, and stochastic noise. This controlled experimental setting enables systematic evaluation of the proposed representation learning strategy without introducing domain-specific confounding effects. Performance is compared against convolutional and patch-based Transformer baselines, including Temporal Convolutional Networks (TCN) [1] and PatchTST [11], and results demonstrate that patch-level temporal tokenization consistently achieves competitive forecasting accuracy on unseen data.
The remainder of this paper is organized as follows. Section II describes the proposed patch-based forecasting framework, detailing the CNN-based patch encoder, token-level self-attention mechanism, and the specialized Transformer architecture used for global temporal modelling. Section III presents the experimental setup, including dataset generation, baseline models, implementation details, and hyperparameter settings. Section IV reports the empirical results and provides a comparative analysis of the proposed method against baseline approaches. Finally, Section V concludes the paper and discusses directions for future research.
@@ II-A Proposed method
Let $\mathbf{X}=\{x_{t}\}_{t=1}^{T}$, where $x_{t}\in\mathbb{R}^{F}$, denote a multivariate time series. Instead of applying a forecasting model directly to $\mathbf{X}$, the proposed method first transforms the input into a sequence of learned temporal tokens. The time series is divided into non-overlapping temporal patches of length $P$, producing $K=\lfloor T/P\rfloor$ segments. Each segment is processed independently using a shared convolutional neural network, which extracts local temporal patterns and non-linear feature interactions. The convolutional outputs are aggregated using temporal pooling and projected into a $D$-dimensional embedding space, yielding a token sequence
$\mathbf{Z}\in\mathbb{R}^{K\times D}$. To incorporate contextual information across segments, multi-head self-attention is applied to the token sequence, enabling interactions across temporal patches during representation learning. The refined token sequence is subsequently used as input to a Transformer encoder, which models inter-patch temporal dependencies and produces patch-level forecasts. This design separates local representation learning from global temporal modelling, leading to more stable and effective forecasting.
@@ II-B CNN Patch Encoder and Attention
Each patch $\mathbf{X}_{k}\in\mathbb{R}^{P\times F}$ is processed independently using a shared one-dimensional convolutional encoder. Convolutions are applied along the temporal dimension, treating the feature dimension as channels, allowing the network to capture short-term temporal patterns and non-linear interactions among features within each patch. Dense connections are employed within the convolutional blocks to preserve intermediate representations and promote feature reuse, enabling both low- and higher-level temporal features to contribute to the final representation.
Following convolutional processing, learned attention-weighted pooling is applied across the temporal dimension of each patch to obtain a fixed-length vector representation. This pooled representation is projected into a $D$-dimensional embedding space, producing an initial token embedding for each patch. To incorporate contextual information across patches, multi-head self-attention is applied to the sequence of token embeddings. This allows each token to attend to others, refining the representations by integrating information from neighbouring and distant temporal segments. The output of this module is a sequence of optimized temporal token embeddings, which serve as structured inputs for subsequent Transformer-based modelling.
@@ III-A2 Seeding
To ensure a fair and reproducible evaluation, two independent datasets were generated using random seeds 42 and 101 for training and testing, respectively. Both datasets follow identical data generation and preprocessing pipelines. There are 10,000 samples in each dataset. All proposed and baseline models are trained exclusively on the training dataset and evaluated on the held-out test dataset.
@@ III-A3 Illustrative Sample
The dynamic features exhibit non-linear and noisy temporal behaviour, including trend and oscillatory components, while the static features act as sequence-level conditioning variables that modulate the target generation process. As shown in the example, static features remain fixed across time steps within a sequence, whereas the dynamic features vary over time. This structure reflects scenarios commonly encountered in financial and economic time-series data, where asset-specific or regime-specific factors remain constant over short horizons while market signals evolve temporally.
@@ IV-A Synthetic Dataset
The observed performance differences can be interpreted in light of the architectural characteristics of the models. PatchTST directly applies self-attention to patch-level representations that are optimized end-to-end for the forecasting objective, which appears advantageous in the considered synthetic setting, particularly under longer input sequences. This tight integration of patch embedding and attention-based modelling enables efficient capture of nonlinear feature interactions.
The TCN baseline exhibits higher error values under the extended sequence length, suggesting limitations in capturing longer-range dependencies when relying solely on convolutional receptive fields. While TCNs effectively model localized temporal structure, their performance degrades relative to attention-based approaches as the temporal context increases.
@@ IV-B Real-World Evaluation: Electricity Load Forecasting
Following standard practice, the data are normalized on a per-meter basis and segmented into sliding windows of fixed length. Each input sample consists of a multivariate sequence of length $T=160$, with the forecasting target defined as the patch-level average load of a selected meter. The same patch size and forecasting horizon used in the synthetic experiments are retained, enabling a consistent evaluation protocol across datasets. Importantly, the convolutional patch encoder trained on synthetic data is reused without architectural modification, highlighting the length-agnostic and transferable nature of the learned local temporal representations.
The proposed model achieves a Mean Squared Error (MSE) of 0.0136 and a Mean Absolute Error (MAE) of 0.0624 on the test split of the Electricity dataset. These results indicate accurate short-horizon forecasting under a real-world setting characterized by strong periodicity, noise, and long-range temporal dependencies. The low MAE value, measured on normalized data, suggests that the model effectively captures both short-term load fluctuations and broader consumption patterns.
The strong performance on electricity load data can be attributed to the architectural separation between local and global modelling stages. Electricity consumption exhibits smooth intra-day dynamics combined with repeating daily and weekly cycles. The convolutional patch encoder operates on short temporal segments, suppressing local noise and extracting stable short-range patterns, while the Transformer focuses on modelling dependencies across patches at a coarser temporal resolution. This division of labour allows the attention mechanism to operate on a more structured and compact representation, which appears particularly effective when longer temporal contexts are considered.
These results demonstrate that the proposed patch-based tokenization strategy generalizes beyond synthetic data and remains effective on a large-scale real-world benchmark. While a detailed comparison with other state-of-the-art electricity forecasting models is beyond the scope of this study, the observed performance provides empirical evidence that decoupling local temporal representation learning from global attention-based modelling yields robust and scalable forecasting behaviour in practical time-series applications.
@@ V Conclusion
This study presents a multi-stage forecasting architecture that explicitly separates patch-level temporal tokenization from global sequence modelling in multivariate time-series forecasting. By learning compact patch-level token embeddings through a convolutional encoder and refining them using token-level self-attention prior to Transformer-based modelling, the proposed framework captures localized temporal structure while enabling effective modelling of longer-range dependencies at a coarser temporal resolution.
Empirical results on a large synthetic dataset with extended sequence length show that the proposed approach consistently outperforms a purely convolutional baseline and achieves competitive performance relative to a strong patch-based Transformer model. Additional evaluation on a real-world electricity load forecasting benchmark demonstrates that the learned patch-level representations transfer effectively beyond controlled synthetic data. While the proposed method does not outperform PatchTST in all settings, it exhibits improved robustness over convolutional baselines and stable performance under increased temporal context. A limitation of this study is the reliance on synthetic data for primary analysis and the use of separate training stages for representation learning and forecasting. Future work will explore jointly trained architectures, alternative patch-level encoders, and broader real-world benchmarks to further assess the generality of the proposed framework.
V-A Limitations and Future Work

### v3

@@ I Introduction
Accurately forecasting multivariate time-series data remains a challenging problem due to the complex, noisy, and often non-stationary nature of many real-world temporal processes. Such time-series data are shaped by a mix of short-lived fluctuations, longer-term trends, and interactions between multiple variables. Together, these factors introduce non-linear behaviour, uneven feature dependencies, and temporal patterns that change over time. As a result, designing forecasting models that perform reliably across different application domains remains a challenging task. Traditional time-series forecasting methods, such as autoregressive and moving-average models [12, 3], are built on assumptions of linearity and stationarity. In many real-world settings, these assumptions do not hold up, which limits the effectiveness of such an approach. This has led to growing interest in deep learning techniques for modelling complex temporal behaviour. Recurrent neural networks (RNNs) [2] have been among the most widely adopted of these methods, as they are specifically designed to process sequential data, particularly Long Short-Term Memory (LSTM) [7] and Gated Recurrent Unit (GRU) [13] architectures, are among the most widely adopted models due to their ability to process sequential data and capture local temporal dependencies. However, recurrent models summarize historical information through compressed hidden states, which can limit their ability to retain long-range dependencies and may lead to unstable training behaviour when long input sequences are considered.
@@
The proposed architecture is initially evaluated on a constructed synthetic dataset designed to reflect key properties and structural dependencies as is observed in real world data, including static and dynamic features as well as stochastic noise. Performance is compared against convolutional and patch-based Transformer baselines: Temporal Convolutional Networks (TCN) [1] and PatchTST [11], and results demonstrate that patch-level temporal tokenization consistently achieves competitive forecasting accuracy on unseen data. Model is then evaluated on a real dataset to show architecture capability in real scenarios and not just tailored situations.
The remainder of this paper is organized as follows. Section II describes the proposed patch-based forecasting framework, detailing the CNN-based patch encoder, token-level self-attention mechanism, and the specialized Transformer architecture used for global temporal modelling. Section III presents the experimental setup, including datasets used and how synthetic dataset was generated, baseline models, implementation details, and hyperparameter settings. Section IV reports the empirical results and provides a comparative analysis of the proposed method against the baseline approaches. Section V concludes the paper and draws subjective conclusions for the proposed model. Section VI discusses the limitations of current architecture, how it can be improved and scope for future works.
@@ II-A Proposed method
Let $\mathbf{X}=\{x_{t}\}_{t=1}^{T}$, where $x_{t}\in\mathbb{R}^{F}$, denote a multivariate time series of length $T$. Rather than applying a forecasting model directly to the raw input sequence, the proposed approach first transforms $\mathbf{X}$ into a sequence of learned temporal tokens. The time series is partitioned into non-overlapping temporal patches of fixed length $P$, resulting in $K=\lfloor T/P\rfloor$ segments. Each patch is processed independently by a shared convolutional neural network, which captures short-range temporal patterns and non-linear interactions among features within the patch. The convolutional representations are subsequently aggregated along the temporal dimension and projected into a $D$-dimensional embedding space, producing a sequence of patch-level tokens $\mathbf{Z}\in\mathbb{R}^{K\times D}$. To enable information exchange across temporal segments, multi-head self-attention is applied to the token sequence during representation learning, allowing dependencies between patches to be explicitly modelled. The refined token sequence is subsequently used as input to a Transformer encoder, which models inter-patch temporal dependencies and produces patch-level forecasts. This design separates local representation learning from global temporal modelling, leading to more stable and effective forecasting.
@@ II-B CNN Patch Encoder and Attention
Each patch $\mathbf{X}_{k}\in\mathbb{R}^{P\times F}$ is processed independently using a shared one-dimensional convolutional encoder. Convolutions are applied along the temporal dimension, treating the feature dimension as channels, allowing the network to capture short-term temporal patterns and non-linear interactions among features within each patch. Dense connections are applied within the convolutional blocks of the encoder to preserve intermediate representations and promote future use, enabling the contribution of both low and higher-level temporal features in final representation. Following the convolutional processing, a learned attention weighted pooling is applied across the temporal dimension of each patch to obtain a fixed length vector representation. This pooled representation is then projected into a $D$-dimensional embedding space, thus producing an initial token embedding for each patch. To incorporate contextual information across patches, multi-head self-attention is applied to the sequence of token embeddings. This allows each token to attend to others, refining the representations by integrating information from neighbouring and distant temporal segments. The output of this module is a sequence of optimized temporal token embeddings, which serve as structured inputs for subsequent Transformer-based modelling.
@@ III-A2 Seeding
To ensure a fair and reproducible evaluation, two independent datasets were generated using random seeds 42 and 101 for training and testing, respectively. Both datasets follow identical data generation and preprocessing pipelines. All proposed and baseline models are trained exclusively on the training dataset and evaluated on the held-out test dataset.
@@ III-A3 Illustrative Sample
The dynamic features exhibit non-linear and noisy temporal behaviour, including trend and oscillatory components, while the static features act as sequence-level conditioning variables that modulate the target generation process. As shown in the dataset snippet, static features are fixed across time-steps within a sequence, the dynamic features, however, vary over time sequence. This structure reflects scenarios commonly encountered in financial and economic time-series data, where asset-specific or regime-specific factors remain constant over short horizons while market signals evolve temporally.
@@ IV-A Synthetic Dataset
The differences in forecasting performance can be understood by considering the architectural design of each model. PatchTST applies self-attention directly to patch-level representations that are learned end-to-end for the forecasting task, which appears beneficial in the synthetic setting considered here, especially when longer input sequences are used. By jointly learning patch embeddings and attention weights, the model is able to capture non-linear feature interactions in an efficient and task-aligned manner.
In contrast, the TCN baseline exhibits higher error under extended sequence lengths, indicating that relying solely on convolutional receptive fields may limit the ability to model longer-range temporal dependencies. While TCNs effectively model localized temporal structure, their performance degrades relative to attention-based approaches as the temporal context increases.
@@ IV-B Real-World Evaluation: Electricity Load Forecasting
Following standard practice, the data are normalized on a per-meter basis and segmented into sliding windows of fixed length. Each input sample consists of a multivariate sequence of length $T=160$, with the forecasting target defined as the patch-level average load of a selected meter. The same patch size and forecasting horizon used in the synthetic experiments are retained to ensure a consistent evaluation protocol across datasets. Notably, the convolutional patch encoder trained on synthetic data is reused without architectural modification, reflecting the length-agnostic nature of the learned local temporal representations.
On the Electricity Load Diagrams dataset, the proposed model achieves a Mean Squared Error (MSE) of 0.0136 and a Mean Absolute Error (MAE) of 0.0624 on the test split. These results indicate accurate short-horizon forecasting in a real-world setting characterized by strong periodic structure, noise, and long-range temporal dependencies. The relatively low MAE value, computed on normalized data, suggests that the model is able to capture both fine-grained load variations and broader consumption trends.
The observed performance can be linked to the architectural separation between local and global modelling stages. Electricity consumption data exhibit smooth intra-day dynamics alongside recurring daily and weekly patterns. The convolutional patch encoder processes short temporal segments, reducing local noise and extracting stable short-range features, while the Transformer focuses on modelling dependencies across patches at a coarser temporal scale. This division of roles enables the attention mechanism to operate on more structured and compact inputs, which appears particularly beneficial when longer temporal contexts are considered.
Overall, these results demonstrate that the proposed patch-based tokenization strategy generalizes beyond synthetic data and remains effective on a large-scale real-world benchmark. Although a detailed comparison with specialized electricity forecasting models is outside the scope of this study, the findings provide empirical support for the benefits of decoupling local temporal representation learning from global attention-based modelling in practical multivariate time-series forecasting tasks.
@@ V Conclusion
This study has aimed to present a multi-stage forecasting model that explicitly separates patch-level tokenization from global sequence modelling in MTSF. By compacting patches through a convolutional encoder and then learning them, refining them through self-attention prior to Transformer modelling, the proposed framework captures localised temporal structure while enabling effective modelling of longer range dependencies.
Empirical results on a large synthetic dataset with extended sequence length show that the proposed approach consistently outperforms a purely convolutional baseline and achieves competitive performance relative to a strong patch-based Transformer model. Additional evaluation on a real-world electricity load forecasting benchmark demonstrates that the learned patch-level representations transfer effectively beyond controlled synthetic data. While the proposed method does not outperform PatchTST in all settings, it exhibits improved robustness over convolutional baselines and stable performance under increased temporal context.
VI Limitations and Future Work